# MESA: Text-Driven Terrain Generation Using Latent Diffusion and Global Copernicus Data

### v1

@@ 1 Introduction
Consequently, a new type of generative model is proposed here and released, referred to as Modeling of Elevation and Surface Appearance (MESA). The model is one of the largest generative models trained on Earth observation data released openly to date, benefiting from global and dense coverage of the open-source Major TOM dataset [5]. Furthermore, to enable this work, a global expansion to Major TOM containing Copernicus DEM in 30-metre resolution has been released under open, free access on HuggingFace (https://huggingface.co/datasets/Major-TOM/Core-DEM). The model architecture based on Stable Diffusion 2.1 produces diverse and realistic terrains represented by corresponding pairs of optical images and elevation maps from text captions. We demonstrate that the model can generate diverse terrain samples of high quality in response to the text prompt, supporting many shapes of terrain, geographical regions and times of the year.
@@ 2 Background
The method described in this work lies at the intersection of several domains (terrain modelling, diffusion models, and remote sensing) and the related works for each of them are summarised below.
@@ 2.1 Terrain Modeling
Current terrain modeling revolves mainly around three types of methods (a detailed overview can be found in [19, 8]): (i) example-based techniques that use real-life terrain as a template to create new ones, (ii) simulations that aim to capture the world complex nature of the world’s landscapes with a finite set of equations and heuristic approximations, and (iii) procedural methods that, without relying on real-life examples or strong physical grounding, focus on producing realistic looking outputs with minimal computational cost and direct control mechanisms. Both simulation and procedural methods commonly suffer from the same pitfall, which can be described as the lack of variability in the type of outputs. This is because each type of terrain (mountain, river, dunes) needs to be modeled explicitly by the method designers and hybrid terrain types are difficult to synthesise.
Furthermore, procedural methods lose realism when the scale is pushed to extremes and simulation computational cost often grows exponentially with the terrain dimensions.
While addressing these two issues, example-based methods were originally limited to patch-based methods yielding suboptimal results when compared to the other methods. More recently machine learning advances especially in the fields of generative modeling spurred a new line of research, for example, terrain authoring using CGANs [14] and more recently, diffusion models [13].
So far these methods have been limited to the authoring of preexisting terrains and do not tackle the problem of general terrain generation. Instead, they enable the user to author preexisting terrains and also are restricted to authoring of elevation maps.
@@ 2.2 Diffusion Models
Diffusion models are generative models that aim to approximate a data distribution $p_{\text{Data}}$ from a finite set of samples [20, 22]. During training, an input image $x\sim p_{\text{Data}}$ is perturbed by adding noise a gaussian noise $\epsilon\sim\mathcal{N}(0,I)$ to create a noisy image $x_{t}=\alpha_{t}x+\sigma_{t}\epsilon$. The noise schedule, parameterized by the diffusion time $t$ (where larger values of $t$ correspond to greater noise levels), is defined by its parameters $\alpha_{t}$ and $\sigma_{t}$
@@ 2.3 Diffusion models for Joint learning of color and depth maps
Physically based rendering (PBR) is a set of techniques in computer graphics that simulates the physical behavior of light to achieve realistic material representation. These approaches rely on material properties maps like roughness, metalness and depth from which the physical properties can be derived.
In our work, we approach PBR material generation as a problem of generating multiple material maps from limited inputs, specifically an albedo map and a depth map. One of the key challenges in this task is handling the increase in output channels. While standard models like Stable Diffusion generate 3-channel RGB outputs, PBR workflows often require outputs with 4, 7, or even 10 channels to account for various material properties such as normal maps, roughness, and metallicity.
Training a new model from scratch, including the VAE and the diffusion model (DM), as proposed in some works, can be computationally expensive and may lead to suboptimal results compared to finetuning existing models. Although finetuning the U-Net component of latent diffusion model (LDM) is relatively straightforward when employing techniques like low-rank adaptation (LoRA) [9] —finetuning the VAE is more complex and tedious. The VAE typically compresses high-frequency information, which remains consistent across different domains (e.g., natural images, satellite images, depth maps), making it viable for encoding and decoding depth maps and other PBR maps
To ensure the diffusion model generates the required material maps in latent space, different strategy can be adopted. We can either assign different prompts to each modality, as done in [23] or modify the architecture to include separate output heads for each modality [12].
Finally, joint learning image appearance (color) and geometry (depth map) in a unified network leads to more realistic outputs by incorporating physical information about the image creation process, as demonstrated in [12].
@@ 3 Method
Building MESA relies on a global dataset of reflectance, elevation and text captions, curated specifically to make this work possible, and a model, based on StableDiffusion, but adapted to the 2.5D representation of terrain.
@@ 3.1.1 Major TOM
Major TOM (Terrestrial Observation Metaset) [5] is an extensible framework for large scale Earth observation dataset insuring machine learning readiness and cross usability of each dataset that follows the framework. The structuring idea behind the dataset is to create a regular grid around the Earth and sample images from modalities of interest for each point. Contrary to regular Earth observation data product format in which images are stored in very large tiles 10 000 pixel wide, samples from Major Tom Core datasets are smaller, and filtered to low cloud coverage, making them immediately applicable in many computer vision deep learning pipelines. The first datasets released within the Major TOM framework already encompass the largest global dataset for Sentinel-1 (S1) and Sentinel-2 (S2) (L2A and L1C). To match this pre-existing dataset we have curated and released a Major TOM dataset for DEM, Major TOM Core-DEM based on Copernicus DEM (GLO30). Using bilinear interpolation the tiles originally in WGS84 were resampled to the UTM projection of the desired grid cell before extracting them, we noticed that choosing bilinear interpolation over nearest neighbor interpolation drastically reduced the stairway artifact issue discussed in [8] while preserving relatively sharp details. The dataset is released under the Major TOM organisation on HuggingFace with free and open access. Both datasets provide global and dense coverage of Earth’s landmass, resulting in a considerable volume of training data indicated in Table 1
@@ 3.1.2 Captions from geographical coordinates
  • A map of countries, following the UN delimitation of borders with this map from the WFP.
@@
  • The topographical information can be more easily captured by describing the Earth’s geology. Using [7] we obtain an indicator that represents the local topology (plain, hills, mountains and so on) and a coarser one that describes the surrounding geological region. This is useful to know the expanse of name entities like "The Alpes" or "The Appalachian". During training, we used both the local description and the regional one.
@@
To get the final extent of the dataset, all cells are considered where it is possible to acquire a corresponding Image and DEM visible as shown on Figure 2. After filtering out oceans and large water bodies with flat elevation, antarctic samples (due to poor signal quality), and cells with misaligned projections or missing data, the training dataset amounts to 1.3 million 10*10km 2.5D terrains which is equivalent to most of Earth’s land surface beyond Antarctica.
@@ 3.2 Terrain Generation
To create our 2.5D terrains we adapt the Stable Diffusion 2.1 architecture to make room for the additional depth channel and finetune the resulting model on the triplet of RGB, DEM, and text. To adjust to the original architecture each dataset image $x_{I}\in\mathbb{R}^{C\times H\times W}$ is cropped to $H=W=768$ and $C=3$. Our DEM cells are notably 3 times coarser than their image counterpart and only require one channel. The initial experiments attempted to take advantage of the coarse spatial resolution of the latent space being close to DEM native resolution and to learn the DEM modality directly as an extra latent channel concatenated to the base. However, this approach did not seem effective and the final design reshapes the DEM so that $x_{D}\in\mathbb{R}^{3\times H\times W}$ and encode it with the same frozen VAE used for $\mathbf{x}_{I}$. This results in two latent vectors $\mathbf{z}_{I}=\mathcal{E}(\mathbf{x}_{I})$ and $,\mathbf{z}_{D}=\mathcal{E}(\mathbf{x}_{D})\in\mathbb{R}^{C^{\prime}\times H^{\prime}\times\ W^{\prime}}$ with $C^{\prime}=4$ and $H^{\prime}=W^{\prime}=96$. Subsequently, the mask is also resized to the dimension of the latent space. The resized mask is represented as $\mathbf{z}_{M}$ for uniformity of notation.
Following an approach similar to previous works [12], the model is trained to simultaneously denoise the depth and the RGB image using two separate modality-specific heads. While the two heads makes the model expressive enough to model the specificity of each modality, the common backbone ensure the spatial alignment and semantic coherence of the modalities (ie. shadows next to hilly areas). We capture the joint distribution of image and DEM with a single U-Net model by simultaneous denoising the different modalities, which can be trained with a simple v-prediction objective. Since the aim is to train a model to generate data without the artifacts due to clouds or missing data, pixels affected by either of these issues are masked out using $\mathbf{z}_{M}$. The final notation is given by
@@
[…]
@@
$t_{I}$, $t_{D}\sim\mathcal{U}[1,T]$ are the sampled time-steps that control the scale of added Gaussian noise. We could use different time-steps for each modality but as mentioned in [12] such strategy only gives $10^{-6}$ probability to sample each perturbation state (given $T=1000$), which may hinder convergence. Therefore we use $t=t_{x}=t_{D}$.
Finally the caption $c\in\mathbb{R}^{L}$ with $L$ being the output dimension of the CLIP text encoder [15] used to compute a vector representation of our caption. Using descriptors introduced in Section 3.1.2 our prompts follow the structure "A Sentinel-2 image of {Biome} and {Geological} in {Country} in {Month}". To ensure that the model does not merely respond to prompts following that exact structure we randomly drop one two or three of these descriptors, we also randomly anonymize the ecoregion name and shuffle between local and regional geological descriptors.
@@
An overview of the architecture can be seen in Figure 3. Using Stable Diffusion 2.1 weights. The frozen image encoders/decoders (VAE) are used to project both RGB and depth maps (concatenated 3 times to match the 3 channel input) to the latent space. During training, both latents are noised (forward diffusion) with independent noise and then jointly denoised conditionaly on caption $c$ (embedded using the frozen text encoder) using a modified U-Net.
@@
The diffusion U-Net architecture consists of down-sampling, middle, and up-sampling blocks, constructed from convolutional layers and self-/cross-attention mechanisms. Specifically, the DownBlocks reduce the resolution of the input noisy latent to generate lower-resolution hidden states, while the UpBlocks perform the reverse, progressively upscaling features to predict the noise. Inspired by [12], we initially considered duplicating the first DownBlock and last UpBlock to create separate heads for each modality, each initialized with Stable Diffusion 2.1 weights. However, we found that using a shared first DownBlock for all modalities yields comparable results while reducing parameter count and simplifying the architecture. The output of this shared head is averaged across modalities then passed to the backbone, whose output is then distributed to the corresponding output heads (last UpBlocks) for each modality.
Using the $\mathbf{z}_{M}$ we mask the loss $\mathcal{L}$ and hence only backpropagate through valid and cloud-free pixels. Assuming that structure is preserved when projecting the input images into the latent space, it enables us to learn a cloud-free distribution of images.
The model is trained on 8 NVIDIA A100 GPUs, with a batch size of 128, fo 80,000 iterations at a constant learning rate of $10^{-5}$ of an AdamW optimizer. For all experiments, we use DDIM sampling with 50 steps and a guidance scale of 7 [21].
@@ 4.2 Influence of captions
Another important experiment to conduct is to test how the model responds to different information provided in captions. MESA can support four types of descriptors in the text prompt: biome, geological, country, and month of the year. Table 3 demonstrates several configurations for information contained in the prompt (rows) for several different samples (columns). The first row contains a prompt with all four descriptors, present (indicated by green icons), while the following rows test for absence of particular pieces of the full prompt. First of all, the date (only used in the first row), is quite important to generate appropriate state of the flora and weather conditions (such as prevent snow from being generated). Biome information is quite important to include to control both features present in the image as well as their scale. The geological has a strong influence over the share of the elevation, and without it, the samples may look quite unrealistic. The exclusion of country information does not have an effect as severe, and as shown in the third last row, including the biome and geological information in the prompt is enough to generate realistic samples. This is also demonstrated in the row, where only country information is provided and the samples look quite unrealistic, leading to very texture-like sample appearance.
@@ 5 Conclusion
This work introduces MESA, a text-driven diffusion model for generating diverse terrains based on features specified using natural language. The model has been trained on a global and dense coverage of Major TOM dataset (minus 10 percent sampled randomly and kept for testing), allowing to account for all kinds of terrain present currently on Earth, indicating a unique value of global and open remote sensing data from the Copernicus programme.
The model is capable of generating 2.5D represenations (optical and depth) of diverse terrains in response to text prompts that control the biome, geological features, country context and season. This has been demonstrated through a series of qualitative experiments and the model exhibits promising performance, to a degree in which it could support creative pipelines where terrain modeling is needed.

### v2

@@ 1 Introduction
Consequently, a new type of generative model is proposed here and released, referred to as Modeling of Elevation and Surface Appearance (MESA). The model is one of the largest generative models trained on Earth observation data released openly to date, benefiting from global and dense coverage of the open-source Major TOM dataset [5]. Furthermore, to enable this work, a global expansion to Major TOM containing Copernicus DEM in 30-metre resolution has been released under open, free access on HuggingFace (https://huggingface.co/datasets/Major-TOM/Core-DEM). The model architecture based on Stable Diffusion 2.1 produces diverse and realistic terrains represented by corresponding pairs of optical images and elevation maps from text captions. We demonstrate that the model can generate diverse terrain samples of high quality in response to the text prompt, supporting many shapes of terrain, geographical regions, and times of the year. The inference code can be found at https://github.com/PaulBorneP/MESA and the model weights at https://huggingface.co/NewtNewt/MESA.
@@ 2 Background
The method described in this work lies at the intersection of several domains (terrain modeling, diffusion models, and remote sensing) and the related works for each of them are summarised below.
@@ 2.1 Terrain Modeling
Current terrain modeling revolves mainly around three types of methods (a detailed overview can be found in [19, 8]): (i) example-based techniques that use real-life terrain as a template to create new ones, (ii) simulations that aim to capture the world complex nature of the world’s landscapes with a finite set of equations and heuristic approximations, and (iii) procedural methods that, without relying on real-life examples or strong physical grounding, focus on producing realistic looking outputs with minimal computational cost and direct control mechanisms. Both simulation and procedural methods commonly suffer from the same pitfall, which can be described as the lack of variability in the type of outputs. This is because each type of terrain (mountain, river, dunes) needs to be modeled explicitly by the 3D Artists, and hybrid terrain types are difficult to synthesise.
Furthermore, procedural methods lose realism when the scale is pushed to extremes, and simulation computational cost often grows exponentially with the terrain dimensions.
While addressing these two issues, example-based methods were originally limited to patch-based algorithms, yielding suboptimal results when compared to the other methods. More recently, machine learning advances, especially in the field of generative modeling, opened a new line of research, for example, terrain authoring using CGANs [14] and, more recently, diffusion models [13].
So far these deep learning methods do not tackle the problem of general terrain generation. Instead, they enable the user to author preexisting terrains and also are restricted to authoring of elevation maps.
@@ 2.2 Diffusion Models
Diffusion models are generative models that aim to approximate a data distribution $p_{\text{Data}}$ from a finite set of samples [20, 22]. During training, an input image $x\sim p_{\text{Data}}$ is perturbed by adding a gaussian noise $\epsilon\sim\mathcal{N}(0,I)$ to create a noisy image $x_{t}=\alpha_{t}x+\sigma_{t}\epsilon$. The noise schedule, parameterized by the diffusion time $t$ (where larger values of $t$ correspond to greater noise levels), is defined by its parameters $\alpha_{t}$ and $\sigma_{t}$
@@ 2.3 Diffusion models for joint learning of color and depth maps
While diffusion models are typically designed to generate standard, 3-channel RGB images, we aim to simultaneously generate both an RGB image and a corresponding depth map using a single unified model. This approach enables the model to learn not only the visual appearance of scenes but also their underlying 3D structure and the relation between the two. Nevertheless this requires a modification to the standard diffusion model. Training a new model from scratch, including the VAE and the diffusion model (DM), as proposed in some works, can be computationally expensive and may lead to suboptimal results compared to finetuning existing models. Although finetuning the U-Net component of latent diffusion model (LDM) is relatively straightforward when employing techniques like low-rank adaptation (LoRA) [9] —finetuning the VAE is more complex and tedious. The VAE typically compresses high-frequency information, which remains consistent across different domains (e.g., natural images, satellite images, depth maps), making it viable for encoding and decoding depth maps.
To ensure that the diffusion model generates the required modalities in latent space, different strategies can be adopted. We can either assign different prompts to each modality, as done in [23] or modify the architecture to include separate output heads for each modality [12].
Finally, jointly learning image appearance (color) and geometry (depth map) in a unified network leads to more realistic outputs by incorporating physical information about the image creation process, as demonstrated in [12].
@@ 3 Method
Building MESA relies on a global dataset of reflectance, elevation, and text captions curated specifically to make this work possible and a model based on StableDiffusion but adapted to the 2.5D representation of terrain.
@@ 3.1.1 Major TOM
Major TOM (Terrestrial Observation Metaset) [5] is an extensible framework for large-scale Earth observation datasets ensuring machine learning readiness and cross usability of each dataset that follows the framework. The structuring idea behind the dataset is to create a regular grid around the Earth and sample images from modalities of interest for each point. Contrary to regular Earth observation data product format in which images are stored in very large tiles 10000 pixels wide, samples from Major Tom Core datasets are smaller and filtered to low cloud coverage, making them immediately applicable in many computer vision deep learning pipelines. The first datasets released within the Major TOM framework already encompass the largest global dataset for Sentinel-1 (S1) and Sentinel-2 (S2) (L2A and L1C). To match this pre-existing dataset, we have curated and released a Major TOM dataset for DEM, Major TOM Core-DEM based on Copernicus DEM (GLO30). Using bilinear interpolation, the tiles originally in WGS84 were resampled to the UTM projection of the desired grid cell before extracting them. We noticed that choosing bilinear interpolation over nearest neighbor interpolation drastically reduced the stairway artifact issue discussed in [8] while preserving relatively sharp details. The dataset is released under the Major TOM organisation on HuggingFace with free and open access. Both datasets provide global and dense coverage of Earth’s landmass, resulting in a considerable volume of training data indicated in Table 1
@@ 3.1.2 Captions from geographical coordinates
  • A map of countries, following the UN (United Nations) delimitation of borders with this map from the WFP (World Food Program).
@@
  • The topographical information can be more easily captured by describing the Earth’s geology. Using [7] we obtain an indicator that represents the local topology (plain, hills, mountains and so on) and a coarser one that describes the surrounding geological region. This is useful to know the expanse of name entities like "The Alps" or "The Appalachian". During training, we used both the local description and the regional one.
@@
To get the final extent of the dataset, all cells are considered where it is possible to acquire a corresponding Image and DEM visible as shown on Figure 2. After filtering out oceans and large water bodies with flat elevation, antarctic samples (due to poor signal quality), and cells with misaligned projections or missing data, the training dataset amounts to 1.3 million 10$\times$10 km2 2.5D terrains which is equivalent to most of Earth’s land surface beyond Antarctica.
@@ 3.2 Terrain Generation
To create our 2.5D terrains, we adapt the Stable Diffusion 2.1 architecture to make room for the additional depth channel and fine-tune the resulting model on the triplet of RGB, DEM, and text. To adjust to the original architecture, each dataset image $x_{I}\in\mathbb{R}^{C\times H\times W}$ is cropped to $H=W=768$ and $C=3$. Our DEM cells are notably 3 times coarser than their image counterpart and only require one channel. The initial experiments attempted to take advantage of the coarse spatial resolution of the latent space being close to DEM native resolution and to learn the DEM modality directly as an extra latent channel concatenated to the base. However, this approach did not seem effective and the final design reshapes the DEM so that $x_{D}\in\mathbb{R}^{3\times H\times W}$ and encode it with the same frozen VAE used for $\mathbf{x}_{I}$. This results in two latent vectors $\mathbf{z}_{I}=\mathcal{E}(\mathbf{x}_{I})$ and $,\mathbf{z}_{D}=\mathcal{E}(\mathbf{x}_{D})\in\mathbb{R}^{C^{\prime}\times H^{\prime}\times\ W^{\prime}}$ with $C^{\prime}=4$ and $H^{\prime}=W^{\prime}=96$. Subsequently, the mask is also resized to the dimension of the latent space. The resized mask is represented as $\mathbf{z}_{M}$ for uniformity of notation.
Following an approach similar to previous works [12], the model is trained to simultaneously denoise the depth and the RGB image using two separate modality-specific heads. While the two heads makes the model expressive enough to model the specificity of each modality, the common backbone ensures the spatial alignment and semantic coherence of the modalities (ie. shadows next to hilly areas). We capture the joint distribution of image and DEM with a single U-Net model by simultaneously denoising the different modalities, which can be trained with a simple v-prediction objective. Since the aim is to train a model to generate data without the artifacts due to clouds or missing data, pixels affected by either of these issues are masked out using $\mathbf{z}_{M}$. The final notation is given by
@@
$\mathbf{v}_{t_{I}}=\alpha_{t_{I}}\epsilon_{I}-\sigma_{t_{I}}\mathbf{z_{I}}$, and $\mathbf{v}_{t_{D}}=\alpha_{t_{D}}\epsilon_{D}-\sigma_{t_{D}}\mathbf{z_{D}}$ the target velocities. Subsequently $\epsilon_{\theta}((\mathbf{v}_{t_{I}},\mathbf{v}_{t_{D}}),c)=(\mathbf{\hat{v}}_{t_{I}},\mathbf{\hat{v}}_{t_{D}})$ with $\mathbf{\hat{v}}_{t_{I}}=\alpha_{t_{I}}\epsilon_{I}-\sigma_{t_{I}}\mathbf{\hat{z}_{I}}$, and $\mathbf{\hat{v}}_{t_{D}}=\alpha_{t_{D}}\epsilon_{D}-\sigma_{t_{D}}\mathbf{\hat{z}_{D}}$ the predicted velocities and $\mathbf{\hat{z}}$ being the result of one step denoising. For ease of notation we write $\mathbf{v}_{I,D,t}=(\mathbf{v}_{t_{I}},\mathbf{v}_{t_{D}})$ and $\mathbf{\hat{v}}_{I,D,t}=(\mathbf{\hat{v}}_{t_{I}},\mathbf{\hat{v}}_{t_{D}})$
@@
$t_{I}$, $t_{D}\sim\mathcal{U}[1,T]$ are the sampled time-steps that control the scale of added Gaussian noise. We could use different time-steps for each modality, but as mentioned in [12], such a strategy only gives $10^{-6}$ probability to sample each perturbation state (given $T=1000$), which may hinder convergence. Therefore we use $t=t_{x}=t_{D}$.
Finally, the caption $c\in\mathbb{R}^{L}$ with $L$ being the output dimension of the CLIP text encoder [15] is used to compute a vector representation of our caption. Using descriptors introduced in Section 3.1.2 our prompts follow the structure "A Sentinel-2 image of {Biome} and {Geological} in {Country} in {Month}". To ensure that the model does not merely respond to prompts following that exact structure we randomly drop one, two, or three of these descriptors. We also randomly anonymize the ecoregion name and shuffle between local and regional geological descriptors.
@@
An overview of the architecture can be seen in Figure 3. Using Stable Diffusion 2.1 weights, the frozen image encoders/decoders (VAE) are used to project both RGB and depth maps (concatenated 3 times to match the 3-channel input) to the latent space. During training, both latents are noised (forward diffusion) with independent noise and then jointly denoised conditionally on caption $c$ (embedded using the frozen text encoder) using a modified U-Net.
@@
The diffusion U-Net architecture consists of down-sampling, middle, and up-sampling blocks, constructed from convolutional layers and self-/cross-attention mechanisms. Specifically, the DownBlocks reduce the resolution of the input noisy latent to generate lower-resolution hidden states, while the UpBlocks perform the reverse, progressively upscaling features to predict the noise. Inspired by [12], we initially considered duplicating the first DownBlock and last UpBlock to create separate heads for each modality, each initialized with Stable Diffusion 2.1 weights. However, we found that using a shared first DownBlock for all modalities yields comparable results while reducing parameter count and simplifying the architecture. The output of this shared head is averaged across modalities and then passed to the backbone, whose output is then distributed to the corresponding output heads (last UpBlocks) for each modality.
Using the $\mathbf{z}_{M}$, we mask the loss $\mathcal{L}$ and hence only backpropagate through valid and cloud-free pixels. Assuming that structure is preserved when projecting the input images into the latent space, it enables us to learn a cloud-free distribution of images.
The model is trained on 8 NVIDIA A100 GPUs, with a batch size of 128, for 80,000 iterations at a constant learning rate of $10^{-5}$ of an AdamW optimizer. For all experiments, we use DDIM sampling with 50 steps and a guidance scale of 7 [21].
@@ 4.2 Influence of captions
Another important experiment to conduct is to test how the model responds to different information provided in captions. MESA can support four types of descriptors in the text prompt: biome, geological, country, and month of the year. Table 3 demonstrates several configurations for information contained in the prompt (rows) for several different samples (columns). The first row contains a prompt with all four descriptors present (indicated by green icons), while the following rows test for the absence of particular pieces of the full prompt. First of all, biome information is quite important to include to control both features present in the image as well as their scale. The geological has a strong influence over the share of the elevation, and without it, the samples may look quite unrealistic. The date (only used in the first row), helps to generate an appropriate state of the flora and weather conditions (such as preventing snow from being generated) but can still be inconsistent in some cases. The exclusion of country information does not have an effect as severe, and as shown in the third last row, including the biome and geological information in the prompt is enough to generate realistic samples. This is also demonstrated in the row where only country information is provided, and the samples look quite unrealistic, leading to a very texture-like sample appearance.
@@ 5 Conclusion
This work introduces MESA, a text-driven diffusion model for generating diverse terrains based on features specified using natural language. The model has been trained on a global and dense coverage of Major TOM dataset (minus 10 percent sampled randomly and kept for testing), allowing to account for all kinds of terrain present currently on Earth. This indicates a unique value of global and open remote sensing data from the Copernicus programme.
The model is capable of generating 2.5D representations (optical and depth) of diverse terrains in response to text prompts that control the biome, geological features, country context, and season. This has been demonstrated through a series of qualitative experiments, and the model exhibits promising performance to a degree in which it could support creative pipelines where terrain modeling is needed.